# Forcing Diffuse Distributions out of Language Models

### v1

@@ 7 Related Work
The lack of diversity has been a long-standing issue in generation (Tevet & Berant, 2021) due to the tension between generation quality and diversity (Zhang et al., 2020): sampling at low temperature causes boring and repetitive text, while sampling at higher temperatures could lead to nonsensical output. Much of the existing literature approaches the problem by coming up with new decoding strategies, which often involve either shaping the model distribution (e.g., top-$p$ sampling (Holtzman et al., 2020) and top-$k$ sampling (Fan et al., 2018)) and diversity-promoting constraints during decoding (Li et al., 2016; Vijayakumar et al., 2018) or training (Welleck et al., 2019; Edunov et al., 2018). Our setting is distinct from prior work in that we already assume a “maximally” diverse decoding strategy (i.e., sampling from the model with temperature one), and language models still fail to produce diverse outputs.
@@
As more capable language models emerge (Touvron et al., 2023; Jiang et al., 2023; Google, 2024), language model-based dataset creation becomes increasingly practical. Prior work in this area has largely focused on generating specialized data for augmenting NLP tasks (Ye et al., 2022b) including semantic similarity (Schick & Schütze, 2021), relationship extraction (Chia et al., 2022), natural language understanding (Meng et al., 2022) and instruction following (Honovich et al., 2022). A LM-based dataset creation pipeline is usually an iterative and arduous process (Ye et al., 2022a), in which significant human intervention is needed to ensure the model generates diverse and high-quality data (Yuan et al., 2021; Liu et al., 2022; Maini et al., 2024). Our work aims to make progress towards the goal of automating data creation by improving the diversity of language model generation and thereby reducing the need for human intervention.
@@ 8 Concluding Discussion and Future Directions
In this work, we propose a method for fine-tuning language models to generate diffuse probability distributions, and show that this method leads to sizable and transferrable improvements in generation diversity. We showcase a practical application of our method in synthetic dataset generation, demonstrating improvements in the quality of generated data by a large margin, with or without task-specific tuning. Our experiments reveal interesting insights on the surprising capability of language models to learn diffuse distributions and generalize to new prompts and output spaces. An important direction for future work is the application of our method (and distribution matching techniques in general) to debiasing language models (Liang et al., 2021), which are shown to be rife with harmful stereotypes (Bolukbasi et al., 2016). Given the strong generalization properties of our method, it is plausible that aligning language models with an ideal distribution on representative instances could lead to generally less biased models. Future work should further explore the limits of this generalization, especially in the context of improving diversity of open-ended generation, where the output spaces are much less structured.
Although our method is independent from the fine-tuning procedure, we find LoRA (Hu et al., 2021) to be substantially more efficient and effective compared to other techniques such as prefix-tuning (Li & Liang, 2021) and discrete prompt search (Zou et al., 2023). Since they could potentially guide closed-source models to generate more diverse outputs, identifying diversity-inducing discrete prompts is a particularly interesting research question, which we leave for future work.
Appendix A Prompts and Qualitative Results
A.1 Prompts used for Figures 2
@@
A.2 Prompt used for synthetic biography generation
@@
A.3 Side-by-side comparison of generated biographies across systems
@@
Appendix B Additional Results
B.1 Results for Baby Names and RNG
Task Model Entropy $\uparrow$ Coverage-1000 $\uparrow$
Baseline Tuned Natural Baseline Tuned Natural
Gemma 1.16 2.28 8 10
RNG Llama-2 1.07 2.29 2.30 7 10 10
Mistral 1.88 2.30 9 10
Gemma 3.19 5.69 170 442
Baby Names Llama-2 3.24 5.83 6.43 129 490 714
Mistral 4.02 5.78 225 479
B.2 Leave-one-out generalization results
@@
B.3 Coverage and entropy of generated biographies
@@
B.4 Maps of birth places in generated biographies
@@
B.5 MT-Bench
We test a version of our models jointly fine-tuned on all tasks in Section 3 on the MT-Bench (Zheng et al., 2023) benchmark. The results show minimal differences between the baseline and our fine-tuned models, suggesting that our optimization does not have a significant impact on the general capabilities of models measured on MT-Bench.

### v2

@@ 7 Related Work
The lack of diversity has been a long-standing issue in generation due to the tension between generation quality and diversity (Zhang et al., 2020): sampling at low temperature causes boring and repetitive text, while sampling at higher temperatures could lead to nonsensical output (Tevet & Berant, 2021). Much of the existing literature approaches the quality-diversity tradeoff by coming up with new decoding strategies, which often involve either shaping the model distribution (e.g., top-$p$ sampling (Holtzman et al., 2020) and top-$k$ sampling (Fan et al., 2018)) and diversity-promoting constraints during decoding (Li et al., 2016; Vijayakumar et al., 2018) or training (Welleck et al., 2019; Edunov et al., 2018). Our setting is distinct from prior work in that we already assume a “maximally” diverse decoding strategy (i.e., sampling from the model with temperature one), and yet instruction-tuned language models still fail to generate diversely.
@@
As more capable language models emerge (Touvron et al., 2023; Jiang et al., 2023; Google, 2024), language model-based dataset creation becomes increasingly practical. Prior work has largely focused on generating specialized data for augmenting NLP tasks (Ye et al., 2022b) including semantic similarity (Schick & Schütze, 2021), relationship extraction (Chia et al., 2022), natural language understanding (Meng et al., 2022) and instruction following (Honovich et al., 2022). A LM-based dataset creation pipeline is usually an iterative and arduous process (Ye et al., 2022a), in which significant human intervention is needed to ensure the model generates diverse and high-quality data (Yuan et al., 2021; Liu et al., 2022; Maini et al., 2024). Our work aims to make progress towards the goal of automating data creation by improving the diversity of language model generation and thereby reducing the need for human intervention.
@@ 8 Concluding Discussion and Future Directions
In this work, we propose a method for fine-tuning language models to generate diffuse probability distributions, and show that this method leads to sizable and transferrable improvements in generation diversity. We showcase a practical application of our method in synthetic dataset generation, demonstrating improvements in the quality of generated data by a large margin, with or without task-specific tuning. Our experiments reveal interesting insights on the surprising capability of language models to learn diffuse distributions and generalize to new prompts and output spaces. An important direction for future work is the application of our method (and distribution matching techniques in general) to debiasing language models (Liang et al., 2021), which are shown to be rife with harmful stereotypes (Bolukbasi et al., 2016). Given the strong generalization properties of our method, it is plausible that aligning language models with an ideal distribution of representative instances could reduce model bias.
Since our method requires access to a set of valid output sequences, a structured output space is assumed. In other words, this method is not directly applicable to open-ended generation tasks where the number of valid targets is practically infinite. Future work should further explore the limits of our approach, especially in the context of improving diversity of open-ended generation. Although our method is independent from the fine-tuning procedure, we find LoRA (Hu et al., 2021) to be substantially more efficient and effective compared to other techniques such as prefix-tuning (Li & Liang, 2021) and discrete prompt search (Zou et al., 2023). Since they could potentially guide closed-source models to generate more diverse outputs, identifying diversity-inducing discrete prompts is a particularly interesting research question, which we leave for future work.
Acknowledgments
We are grateful to the Center For AI Safety for providing compute for experiments in this paper.
Appendix A Training details
For all experiments, we use LoRA to fine-tune models for up to 50 steps over all target prompts. In tasks where the ground truth distribution cannot be enumerated (e.g., synthetic biographies), we stop fine-tuning early when the training loss is within $20\%$ of the minimum loss, which is the entropy of the target distribution. We use an AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $5\times 10^{-5}$ and a batch size of 32. All experiments are conducted on a single NVIDIA A100 GPU, using a total of approximately 50 GPU hours. Code and prompts used to reproduce our results are available at the open-source GitHub repository. [^13] https://github.com/y0mingzhang/diffuse-distributions
Appendix B Prompts and Qualitative Results
B.1 Prompts used for Figures 2
@@
B.2 Prompt used for synthetic biography generation
@@
B.3 Side-by-side comparison of generated biographies across systems
@@
Appendix C Additional Results
C.1 Leave-one-out generalization results
@@
C.2 Coverage and entropy of generated biographies
@@
C.3 Maps of birth places in generated biographies
@@
C.4 MT-Bench
After fine-tuning the models on all six tasks in Section 5, we evaluate utility of the models on the MT-Bench (Zheng et al., 2023) benchmark. The results show minimal differences between the baseline and our fine-tuned models, suggesting that our optimization does not have a significant impact on the general capabilities of models.